# SMReferences for Appendix

### v3

@@ 3 Average treatment effects: theoretical analysis
In this section, we derive the asymptotic result for the proposed DCal estimator $\hat{\tau}_{\textrm{DCal}}$. For simplicity of exposition, throughout this section and the next section, we consider an asymptotic regime where the dimension $p$ grows with $n$ and $n\equiv n_{\mathrm{aux}}\equiv n_{\mathrm{tr}}$. We first prove that $\hat{\tau}_{\textrm{DCal}}$ is $\sqrt{n}$-consistent when the OR follows a sparse GLM model in Section 3.1, followed by a proof of its $\sqrt{n}$-consistency when the PS follows a sparse GLM model in Section 3.2.
@@ 3.1 Theoretical properties under the sparse OR model
Recall that our target of interest is the ATE over the main sample $\bar{\tau}^{\ast}$ in (3). In this section, we prove the $\sqrt{n}$-consistency of $\hat{\tau}_{\textrm{DCal}}$ for estimating $\bar{\tau}^{\ast}$ when only the OR model is sparse. Throughout Section 3, we assume the following regularity conditions.
Condition 4.
For some constant $m_{Y}>0$, almost surely, $|Y(1)|,|r^{*}(X)|\leq m_{Y}$.
We also invoke the following standard condition on the design matrix $\mathbf{X}$:
Condition 5.
The first component of $X$ is $1$, representing an intercept term. Denoting by $Z\in\mathbb{R}^{p-1}$ the remaining components of $X$, we assume $\mathbb{E}Z=0$ and almost surely, $\|Z\|_{\infty}\leq m_{Z}$. Moreover, $\mathbf{X}$ is full rank and there exists a constant $\sigma_{Z}$ such that for any unit vector $u\in\mathbb{R}^{p-1}$ and all $\alpha\in\mathbb{R}$, $\mathbb{E}\{\exp(\alpha u^{\top}Z)\}\leq\exp(\alpha^{2}\sigma_{Z}^{2}/2)$.
As we have mentioned in the comments after Lemma 1, we say that “$\mathbf{X}$ is full rank” if the rank of $\mathbf{X}$ equals $\min\{n,p\}$. In addition, we need the following conditions on the nuisance parameter estimates computed from the training and the auxiliary datasets. Recall that $\hat{\pi}_{\mathrm{aux}}$ and $\hat{r}_{\mathrm{aux}}$ appears in the constraint (5) when calibrating the OR model.
@@
There exists a constant $m_{r}>0$ such that with probability converging to $1$, for all $i$,
$|\hat{r}_{i}|\leq m_{r}\quad\&\quad|\hat{r}_{\mathrm{aux},i}|\leq m_{r}.$
Condition 7.
@@
We then have the following important observation (Lemma 2): when the initial OR estimate $\hat{r}$ is close to the true OR $r^{\ast}$, correspondingly $\|\hat{\bm{\mu}}\|_{2}$ is small. This observation tells us that the OR-calibration is harmless even if it is superfluous when $\hat{r}$ is already close to $r^{\ast}$. This is appealing because in practice, we generally have no idea if $\hat{r}$ converges to $r^{\ast}$ at a sufficiently fast rate even under the (minimal sparsity) condition.
@@
Suppose Conditions 12 and 47 hold with $s_{r}=o(\sqrt{n}/\log p)$, then by choosing $\eta_{r}\asymp\sqrt{\frac{\log p}{n}}$, $M_{r},M_{\gamma}\asymp 1$ sufficiently large, we have that $\|\hat{\bm{\mu}}\|_{2}=o_{\mathbb{P}}(n^{1/4})$.
The claim of Lemma 2 is expected as $\hat{\bm{\mu}}$, by design, calibrates the estimation error $r_{i}^{*}-\hat{r}_{i}$. When $\hat{r}_{i}$ estimates $r_{i}^{*}$ with high accuracy, intuitively one should expect the calibrator $\hat{\bm{\mu}}$ to be negligible. Now equipped with the above result, using Taylor expansion, we can decompose $\hat{\tau}_{\textrm{DCal}}-\bar{\tau}^{\ast}$ as
$\hat{\tau}_{\textrm{DCal}}$ $-\bar{\tau}^{\ast}$
@@
where $\varepsilon_{i}(1)\coloneqq Y_{i}(1)-r^{*}(X_{i})$ is the residual noise of the random variable $Y_{i}(1)$. As explained in Remark 2, by design of the program (8), the last term is mean zero and asymptotically normal. Therefore the first two terms constitute the bias. By applying Hölder’s inequality and using the constraint of $\tilde{\pi}_{i}$ in (9) and (10), we have that the first two terms are, with probability converging to one, of order
@@
Suppose Conditions 12 and 47 hold with $s_{r}=o(\sqrt{n}/\log p)$, by choosing $\eta_{r},\eta_{\pi_{1}}\asymp\sqrt{\frac{\log p}{n}},\eta_{\pi_{2}}\asymp\sqrt{\frac{\log p\vee n}{n}},M_{r},M_{\pi},M_{\gamma}\asymp 1$ sufficiently large, we have the following representation
$\sqrt{n}(\hat{\tau}_{\textrm{DCal}}-\bar{\tau}^{\ast})=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\frac{T_{i}\varepsilon_{i}(1)}{\tilde{\pi}_{i}}+o_{\mathbb{P}}(1).$
Let $\bar{\sigma}_{r}^{2}\coloneqq n^{-1}\sum_{i=1}^{n}T_{i}\textrm{var}(\varepsilon_{i}(1)\mid X_{i})/\tilde{\pi}_{i}^{2}$. Informally, Theorem 3 states that with sufficiently large $n$, with high probability, we can obtain by conditioning on $\{(\mathbf{X},\mathbf{T}),\mathcal{D}_{\mathrm{aux}},\mathcal{D}_{\mathrm{tr}}\}$ that
$\sqrt{n}(\hat{\tau}_{\textrm{DCal}}-\bar{\tau}^{\ast})\mid\{(\mathbf{X},\mathbf{T}),\mathcal{D}_{\mathrm{aux}},\mathcal{D}_{\mathrm{tr}}\}\approx\mathcal{N}(0,\bar{\sigma}_{r}^{2}).$
Once we further have the assumption that almost surely, $\textrm{var}(\varepsilon_{i}(1)\mid X_{i})$ is bounded above by some constant, we have that with probability converging to $1$, $\bar{\sigma}_{r}$ is $O(1)$. This means that $\hat{\tau}_{\textrm{DCal}}$ is a $\sqrt{n}$-consistent estimator for $\bar{\tau}^{\ast}$. In fact, our convergence guarantee is quite similar to the one given by Athey et al. (2018), except that they replaced $1/\tilde{\pi}_{i}$ by some weights learned by a convex program.
@@
We conjecture that it should be the case under the additional assumption that $\pi^{\ast}$ is consistently estimable by $\tilde{\pi}$ when only the OR is assumed to be sparse. ∎
@@ 3.2 Theoretical properties under the sparse PS model
In this section, we consider the other way around, namely when the PS model satisfies a sparse GLM. Specifically, we impose the following assumption on the estimated PS coefficient $\hat{\gamma}$ computed from the training sample $\mathcal{D}_{\mathrm{tr}}$:
Condition 8.
@@
This, together with Condition 2, implies that with probability converging $1$, by choosing $M_{\gamma}\asymp 1$ large enough, $|X_{i}^{\top}\hat{\gamma}|,|X_{\mathrm{aux},i}^{\top}\hat{\gamma}|\leq M_{\gamma}$, i.e., that with probability converging to $1$, $\hat{\pi}_{i}\equiv\phi(X_{i}^{\top}\hat{\gamma})$ and $\hat{\pi}_{\mathrm{aux},i}\equiv\phi(X_{\mathrm{aux},i}^{\top}\hat{\gamma})$. This is helpful to derive the $\sqrt{n}$-consistency of $\hat{\tau}_{\textrm{DCal}}$.
We also impose the following constraint on the “lower bound” of $\phi^{\prime}$:
Condition 9.
@@
Condition 9 is satisfied by e.g. logistic and probit regression models. With these additional constraints on the sparsity and link functions, the calibrated $\tilde{\gamma}$ satisfies that
@@
Suppose Conditions 1246 and Conditions 89 hold; suppose moreover that $s_{\pi}=o(\sqrt{n}/\log p)$. By choosing the tuning parameters as in Theorem 3, we have that with probability converging to $1$,
@@
Armed with this lemma, and recall that we already have $\hat{\tau}_{\textrm{SCal},r}$ is $\sqrt{n}$-consistent, it remains to analyze the difference between $\hat{\tau}_{\textrm{DCal}}$ and $\hat{\tau}_{\textrm{SCal},r}$. We now apply the following decomposition
$\begin{split}\hat{\tau}_{\textrm{DCal}}-\hat{\tau}_{\textrm{SCal},r}=&\ \frac{1}{n}\sum_{i=1}^{n}\left(\frac{T_{i}(Y_{i}-\hat{r}_{i}-\hat{\mu}_{i})}{\tilde{\pi}_{i}}-\frac{T_{i}(Y_{i}-\hat{r}_{i}-\hat{\mu}_{i})}{\hat{\pi}_{i}}\right)\\ =&\ \frac{1}{n}\sum_{i=1}^{n}\left(\frac{T_{i}(Y_{i}-\hat{r}_{i})}{\tilde{\pi}_{i}\hat{\pi}_{i}}-\frac{\hat{\mu}_{i}}{\hat{\pi}_{i}}\right)(\hat{\pi}_{i}-\tilde{\pi}_{i})+\frac{1}{n}\sum_{i=1}^{n}(T_{i}-\tilde{\pi}_{i})\frac{\hat{\mu}_{i}}{\hat{\pi}_{i}\tilde{\pi}_{i}}(\hat{\pi}_{i}-\tilde{\pi}_{i}).\end{split}$
@@
$\begin{split}\hat{\tau}_{\textrm{DCal}}-\hat{\tau}_{\textrm{SCal},r}\lesssim&\ \left|\frac{1}{n}\sum_{i=1}^{n}\left(\frac{T_{i}(Y_{i}-\hat{r}_{i})}{\tilde{\pi}_{i}\hat{\pi}_{i}}-\frac{\hat{\mu}_{i}}{\hat{\pi}_{i}}\right)\phi^{\prime}(X_{i}^{\top}\hat{\gamma})\tilde{X}_{i}^{\top}(\tilde{\gamma}-\hat{\gamma})\right|\\ &+\left|\frac{1}{n}\sum_{i=1}^{n}\left(\frac{T_{i}}{\tilde{\pi}_{i}}-1\right)\frac{\phi^{\prime}(X_{i}^{\top}\hat{\gamma})\hat{\mu}_{i}}{\hat{\pi}_{i}}\tilde{X}_{i}^{\top}(\tilde{\gamma}-\hat{\gamma})\right|+\frac{1}{n}\sum_{i=1}^{n}(X_{i}^{\top}(\tilde{\gamma}-\gamma^{*}))^{2}.\end{split}$ (14)
@@
Suppose Conditions 1246 and Conditions 89 hold with $s_{\pi}=o(\sqrt{n}/\sqrt{\log p\log p\vee n})$, then by choosing we the tuning parameters in the same way as in Theorem 3, we have the following representation
$\sqrt{n}(\hat{\tau}_{\textrm{DCal}}-\bar{\tau}^{\ast})=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left(\frac{T_{i}\varepsilon_{i}(1)}{\pi_{i}^{*}}+\frac{(T_{i}-\pi_{i}^{*})(r_{i}^{*}-\tilde{r}_{i})}{\pi_{i}^{*}}\right)+o_{\mathbb{P}}(1).$
@@
$\sigma_{\mu}^{2}\coloneqq\frac{1}{n}\sum_{i=1}^{n}\frac{(1-\pi_{i}^{*})(r_{i}^{*}-\tilde{r}_{i})^{2}}{\pi_{i}^{*}},\quad\bar{\sigma}_{\pi}^{2}\coloneqq\frac{1}{n}\sum_{i=1}^{n}\frac{\textrm{var}(\varepsilon_{i}(1)^{2}\mid X_{i})}{\pi_{i}^{*}}.$
@@
$\sqrt{n}(\hat{\tau}_{\textrm{DCal}}-\bar{\tau}^{\ast})\mid\{\mathbf{X},\mathcal{D}_{\mathrm{aux}},\mathcal{D}_{\mathrm{tr}}\}\approx\mathcal{N}(0,\bar{\sigma}_{\pi}^{2}+\sigma_{\mu}^{2}).$
As also discussed in Theorem 3, under some standard regularity conditions, one can expect $\bar{\sigma}_{\pi}$ to be $O(1)$. For $\sigma_{\mu}^{2}$, using the overlap condition as in Condition 2 and recall the $\ell_{\infty}$ constraint (6), one can expect $\sigma^{2}_{\mu}$ to be of order $O(1)$. This proves the $\sqrt{n}$-consistency of our estimator under a sparse PS model. This asymptotic distribution is similar to the one in Wang and Shah (2020).
@@
Similar to the comment in Remark 3, one missing piece of Theorem 5 is if the asymptotic linear representation can achieve the semiparametric efficiency bound (with respect to the main dataset $\mathcal{D}$). We again conjecture that it should be the case under the additional assumption that $r^{\ast}$ is consistently estimable by $\tilde{r}$ when only the PS is assumed to be sparse. ∎
To conclude, Theorem 3 and 5, taken together, address Problem* when $p>n$ as the sparsity constraint $s_{\pi}\ll\sqrt{n}/\sqrt{\log p\log p\vee n}$ reduces to $s_{\pi}\ll\sqrt{n}/\log p$. When $p\leq n$, the DCal estimator achieves $\sqrt{n}$-consistency under the (minimal sparsity) condition up to a minor $\sqrt{\log n/\log p}$ factor, which is a constant if we take $p=n^{\delta}$ for some fixed $\delta\in(0,1)$. We conjecture that this extra log factor can be removed by calibrating the PS in a similar fashion to the OR calibration program (4), which we leave to future work.
@@
The current analysis still requires that the observed data $O=(X,T,Y)$ be light-tailed. It is of great interest to investigate if the proposed approach can be extended into settings in which (part of) $O$ has heavier tails (Kuchibhotla and Chakrabortty, 2022) or has even diverging higher-order moments (e.g. transformed covariates by B-splines or Cohen-Daubechies-Vial wavelets (Cohen et al., 1993; Belloni et al., 2015; Liu et al., 2017; Mukherjee and Sen, 2018; Liu et al., 2021)) [^8] Based on personal communications (Mukherjee, 2023), R. Mukherjee and collaborators have established nontrivial (and possibly tight) lower bounds for certain doubly-robust functionals in this context in an unpublished technical report.. ∎
4 Some extensions of the Double-Calibration approach
In this section, we extend the Double-Calibration approach for ATE estimation in two directions.

### v4

@@ 3 Average treatment effects: theoretical analysis
In this section, we derive the asymptotic result for the proposed DCal estimator $\hat{\tau}_{\mathrm{DCal}}$. For simplicity, throughout this section and the next section, we take $n\equiv n_{\mathrm{aux}}\equiv n_{\mathrm{tr}}$ and assume throughout that we are under an asymptotic regime where $p\to\infty$ as $n\to\infty$ and that $\log p/\sqrt{n}=o(1)$. We also assume throughout the following conditions on the design matrix and outcomes as well as the estimated outcome regression model. Recall that $\hat{\pi}_{\mathrm{aux}}$ and $\hat{r}_{\mathrm{aux}}$ appears in the constraint (6b) when calibrating the OR model.
Condition 4.
We have that
  • (i)
    For some constant $m_{Y}>0$, almost surely, $|Y(1)|,|r^{*}(X)|\leq m_{Y}$;
  • (ii)
    The first component of $X$ is $1$, representing an intercept term. Denoting by $V\in\mathbb{R}^{p-1}$ the remaining components of $X$, we assume $\mathbb{E}V=0$ and almost surely, $\|V\|_{\infty}\leq m_{V}$. Moreover, $\mathbf{X}$ is full rank and there exists a constant $\sigma_{V}$ such that for any unit vector $u\in\mathbb{R}^{p-1}$ and all $\alpha\in\mathbb{R}$, $\mathbb{E}\{\exp(\alpha u^{\top}V)\}\leq\exp(\alpha^{2}\sigma_{V}^{2}/2)$.
Condition 5.
There exists a constant $m_{r}>0$ such that with probability converging to $1$, $\mathbb{P}(|\psi(X^{\top}\hat{\beta})|\leq m_{r}\mid\hat{\beta})=1$, where $X$ is a random vector that is independent from $\hat{\beta}$ and follows the same distribution as the $X_{i}$’s.
Notice that as we have mentioned in the comments after Lemma 1, we say that “$\mathbf{X}$ is full rank” if the rank of $\mathbf{X}$ equals $\min\{n,p\}$. Notice also that the $\ell_{\infty}$ constraint on $V$ can be removed by imposing additional constraints in the optimization programs (6a) and (9a). The purpose of the boundedness assumption imposed on $Y$ is for us to focus on the main thesis of the paper – constructing a $\sqrt{n}$-consistent estimator of $\tau$ under (minimal sparsity). We expect that similar conclusions should hold under light-tail assumptions such as sub-Gaussianity. In the rest of this section, we first prove that $\hat{\tau}_{\mathrm{DCal}}$ is $\sqrt{n}$-consistent when the OR follows a sparse GLM model in Section 3.1, followed by a proof of its $\sqrt{n}$-consistency when the PS follows a sparse GLM model in Section 3.2.
@@ 3.1 Theoretical properties under the sparse OR model
Recall that our target of interest is the ATE over the main sample $\bar{\tau}^{\ast}$ in (4). In this section, we prove the $\sqrt{n}$-consistency of $\hat{\tau}_{\mathrm{DCal}}$ for estimating $\bar{\tau}^{\ast}$ when only the OR model is sparse. To achieve this goal, we invoke the following conditions on the nuisance parameter estimates computed from the training and the auxiliary datasets.
@@
Condition 6 can be met by using common $\ell_{1}$-regularized techniques such as the lasso (Tibshirani, 1996), square-root lasso (Belloni et al., 2011), SLOPE (Su and Candes, 2016; Bellec et al., 2018) when the penalized loss function satisfies certain regularity conditions such as strong convexity.
We then have the following important observation (Lemma 2): when the initial OR estimate $\hat{r}$ is close to the true OR $r^{\ast}$, correspondingly $\|\hat{\bm{\mu}}\|_{2}$ is small. This observation tells us that the OR-calibration is harmless when $\hat{r}$ is already close to $r^{\ast}$. This is appealing because in practice, we generally have no idea if $\hat{r}$ converges to $r^{\ast}$ at a sufficiently fast rate even under the (minimal sparsity) condition.
@@
Suppose Conditions 12 and 46 hold with $s_{r}=o(\sqrt{n}/\log p)$ and let the constant $M_{\gamma}$ be given; then by choosing $\eta_{r}\asymp\sqrt{\frac{\log p}{n}}$, $M_{r},M_{\gamma}\asymp 1$ sufficiently large, $\|\hat{\bm{\mu}}\|_{2}=o_{\mathbb{P}}(n^{1/4})$.
The claim of Lemma 2 is expected as $\hat{\bm{\mu}}$, by design, calibrates the estimation error $r_{i}^{*}-\hat{r}_{i}$. When $\hat{r}_{i}$ estimates $r_{i}^{*}$ with high accuracy, intuitively one should expect the calibrator $\hat{\bm{\mu}}$ to be negligible. Now equipped with the above result, using Taylor expansion, we can decompose $\hat{\tau}_{\mathrm{DCal}}-\bar{\tau}^{\ast}$ as
$\hat{\tau}_{\mathrm{DCal}}$ $-\bar{\tau}^{\ast}$
@@
where $\varepsilon_{i}(1)\coloneqq Y_{i}(1)-r^{*}(X_{i})$ is the residual noise of the random variable $Y_{i}(1)$. As explained in Remark 2, by design of the program (9a), the last term is mean zero and asymptotically normal. Therefore the first two terms constitute the bias. By applying Hölder’s inequality and using the constraint of $\tilde{\pi}_{i}$ in (9b) and (9c), we have that the first two terms are, with probability converging to one, of order
@@
Suppose Conditions 12 and 46 hold with $s_{r}=o(\sqrt{n}/\log p)$ and let the constant $M_{\gamma}$ be given, by choosing $\eta_{r},\eta_{\pi_{1}}\asymp\sqrt{\frac{\log p}{n}},\eta_{\pi_{2}}\asymp\sqrt{\frac{\log p\vee n}{n}},M_{r},M_{\pi}\asymp 1$ sufficiently large, we have the following representation
$\sqrt{n}(\hat{\tau}_{\mathrm{DCal}}-\bar{\tau}^{\ast})=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\frac{T_{i}\varepsilon_{i}(1)}{\tilde{\pi}_{i}}+o_{\mathbb{P}}(1).$
Let $\bar{\sigma}_{r}^{2}\coloneqq n^{-1}\sum_{i=1}^{n}T_{i}\mathrm{var}(\varepsilon_{i}(1)\mid X_{i})/\tilde{\pi}_{i}^{2}$. Informally, Theorem 3 states that with sufficiently large $n$, with high probability, we can obtain by conditioning on $\{(\mathbf{X},\mathbf{T}),\mathcal{D}_{\mathrm{aux}},\mathcal{D}_{\mathrm{tr}}\}$ that
$\sqrt{n}(\hat{\tau}_{\mathrm{DCal}}-\bar{\tau}^{\ast})\mid\{(\mathbf{X},\mathbf{T}),\mathcal{D}_{\mathrm{aux}},\mathcal{D}_{\mathrm{tr}}\}\approx\mathcal{N}(0,\bar{\sigma}_{r}^{2}).$
Once we further have the assumption that almost surely, $\mathrm{var}(\varepsilon_{i}(1)\mid X_{i})$ is bounded above by some constant, we have that with probability converging to $1$, $\bar{\sigma}_{r}$ is $O(1)$. This means that $\hat{\tau}_{\mathrm{DCal}}$ is a $\sqrt{n}$-consistent estimator for $\bar{\tau}^{\ast}$. In fact, our convergence guarantee is quite similar to the one given by Athey et al. (2018), except that they replaced $1/\tilde{\pi}_{i}$ by some weights learned by solving a convex program.
@@
We conjecture that it should be the case under the additional assumption that $\pi^{\ast}$ can be consistently estimated by $\tilde{\pi}$ when only the OR is assumed to be sparse. ∎
@@ 3.2 Theoretical properties under the sparse PS model
In this section, we consider “the other way around”, namely when the PS model satisfies a sparse GLM. Specifically, we impose the following assumption on the estimated PS coefficient $\hat{\gamma}$ computed from the training sample $\mathcal{D}_{\mathrm{tr}}$:
Condition 7.
@@
This, together with Condition 2, implies that with probability converging $1$, by choosing $M_{\gamma}\asymp 1$ large enough, $|X_{i}^{\top}\hat{\gamma}|,|X_{\mathrm{aux},i}^{\top}\hat{\gamma}|\leq M_{\gamma}$, i.e., that with probability converging to $1$, $\hat{\pi}_{i}\equiv\phi(X_{i}^{\top}\hat{\gamma})$ and $\hat{\pi}_{\mathrm{aux},i}\equiv\phi(X_{\mathrm{aux},i}^{\top}\hat{\gamma})$. This is helpful to derive the $\sqrt{n}$-consistency of $\hat{\tau}_{\mathrm{DCal}}$.
We also impose the following “lower bound” constraint on $\phi^{\prime}$:
Condition 8.
@@
Condition 8 is satisfied by e.g. logistic and probit regression models. With these additional constraints on the sparsity and link functions, the calibrated $\tilde{\gamma}$ satisfies that
@@
Suppose Conditions 1245 and Conditions 78 hold; suppose moreover that $s_{\pi}=o(\sqrt{n}/\log p)$. By choosing $\eta_{r},\eta_{\pi_{1}}\asymp\sqrt{\frac{\log p}{n}},\eta_{\pi_{2}}\asymp\sqrt{\frac{\log p\vee n}{n}},M_{r},M_{\pi},M_{\gamma}\asymp 1$ sufficiently large, we have that with probability converging to $1$,
@@
Armed with this lemma, and recall that $\hat{\tau}_{\textrm{SCal},r}$ is $\sqrt{n}$-consistent (see Appendix A.5 for an elaboration of the analysis), it remains to analyze the difference between $\hat{\tau}_{\mathrm{DCal}}$ and $\hat{\tau}_{\textrm{SCal},r}$. We now apply the following decomposition
$\begin{split}\hat{\tau}_{\mathrm{DCal}}-\hat{\tau}_{\textrm{SCal},r}=&\ \frac{1}{n}\sum_{i=1}^{n}\left(\frac{T_{i}(Y_{i}-\hat{r}_{i}-\hat{\mu}_{i})}{\tilde{\pi}_{i}}-\frac{T_{i}(Y_{i}-\hat{r}_{i}-\hat{\mu}_{i})}{\hat{\pi}_{i}}\right)\\ =&\ \frac{1}{n}\sum_{i=1}^{n}\left(\frac{T_{i}(Y_{i}-\hat{r}_{i})}{\tilde{\pi}_{i}\hat{\pi}_{i}}-\frac{\hat{\mu}_{i}}{\hat{\pi}_{i}}\right)(\hat{\pi}_{i}-\tilde{\pi}_{i})+\frac{1}{n}\sum_{i=1}^{n}(T_{i}-\tilde{\pi}_{i})\frac{\hat{\mu}_{i}}{\hat{\pi}_{i}\tilde{\pi}_{i}}(\hat{\pi}_{i}-\tilde{\pi}_{i}).\end{split}$
@@
$\begin{split}\hat{\tau}_{\mathrm{DCal}}-\hat{\tau}_{\textrm{SCal},r}\lesssim&\ \left|\frac{1}{n}\sum_{i=1}^{n}\left(\frac{T_{i}(Y_{i}-\hat{r}_{i})}{\tilde{\pi}_{i}\hat{\pi}_{i}}-\frac{\hat{\mu}_{i}}{\hat{\pi}_{i}}\right)\phi^{\prime}(X_{i}^{\top}\hat{\gamma})\tilde{X}_{i}^{\top}(\tilde{\gamma}-\hat{\gamma})\right|\\ &+\left|\frac{1}{n}\sum_{i=1}^{n}\left(\frac{T_{i}}{\tilde{\pi}_{i}}-1\right)\frac{\phi^{\prime}(X_{i}^{\top}\hat{\gamma})\hat{\mu}_{i}}{\hat{\pi}_{i}}\tilde{X}_{i}^{\top}(\tilde{\gamma}-\hat{\gamma})\right|+\frac{1}{n}\sum_{i=1}^{n}(X_{i}^{\top}(\tilde{\gamma}-\gamma^{*}))^{2}.\end{split}$ (12)
@@
Suppose Conditions 1245 and Conditions 78 hold with $s_{\pi}=o(\sqrt{n}/\sqrt{\log p\log p\vee n})$, then by choosing we the tuning parameters in the same way as in Theorem 3, we have the following representation
$\sqrt{n}(\hat{\tau}_{\mathrm{DCal}}-\bar{\tau}^{\ast})=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left(\frac{T_{i}\varepsilon_{i}(1)}{\pi_{i}^{*}}+\frac{(T_{i}-\pi_{i}^{*})(r_{i}^{*}-\tilde{r}_{i})}{\pi_{i}^{*}}\right)+o_{\mathbb{P}}(1).$
@@
$\sigma_{\mu}^{2}\coloneqq\frac{1}{n}\sum_{i=1}^{n}\frac{(1-\pi_{i}^{*})(r_{i}^{*}-\tilde{r}_{i})^{2}}{\pi_{i}^{*}},\quad\bar{\sigma}_{\pi}^{2}\coloneqq\frac{1}{n}\sum_{i=1}^{n}\frac{\mathrm{var}(\varepsilon_{i}(1)^{2}\mid X_{i})}{\pi_{i}^{*}}.$
@@
$\sqrt{n}(\hat{\tau}_{\mathrm{DCal}}-\bar{\tau}^{\ast})\mid\{\mathbf{X},\mathcal{D}_{\mathrm{aux}},\mathcal{D}_{\mathrm{tr}}\}\approx\mathcal{N}(0,\bar{\sigma}_{\pi}^{2}+\sigma_{\mu}^{2}).$
As also discussed in Theorem 3, under some standard regularity conditions, one can expect $\bar{\sigma}_{\pi}$ to be $O(1)$. For $\sigma_{\mu}^{2}$, using the overlap condition as in Condition 2 and recall the $\ell_{\infty}$ constraint (6c), one can expect $\sigma^{2}_{\mu}$ to be of order $O(1)$. This proves the $\sqrt{n}$-consistency of our estimator under a sparse PS model. This asymptotic distribution is similar to the one in Wang and Shah (2024).
@@
Similar to the comment in Remark 3, one missing piece of Theorem 5 is if the asymptotic linear representation can achieve the semiparametric efficiency bound (with respect to the main dataset $\mathcal{D}$). We again conjecture that it should be the case under the additional assumption that $r^{\ast}$ can be consistently estimated by $\tilde{r}$ when only the PS is assumed to be sparse. ∎
To conclude, Theorem 3 and 5, taken together, address Problem* when $p>n$ as the sparsity constraint $s_{\pi}\ll\sqrt{n}/\sqrt{\log p\log p\vee n}$ reduces to $s_{\pi}\ll\sqrt{n}/\log p$. When $p\leq n$, the DCal estimator achieves $\sqrt{n}$-consistency under the (minimal sparsity) condition up to a minor $\sqrt{\log n/\log p}$ factor, which is a constant if we take $p=n^{\delta}$ for some fixed $\delta\in(0,1)$. We conjecture that this extra log factor can be removed by calibrating the PS in a similar fashion to the OR calibration program (6a), which we leave to future work.
We now briefly discuss the necessity of (minimal sparsity) for $\sqrt{n}$-consistency of the ATE. In terms of the matching minimax lower bounds, to the best of our knowledge, Bradic et al. (2019a) was the first to show that (minimal sparsity) is necessary for $\sqrt{n}$-consistency for continuous $T$ when both the PS and OR models are linear. As mentioned in the Introduction, they also constructed two separate $\sqrt{n}$-consistent estimators when the PS or the OR is sufficiently sparse. At referees’ request during the revision process, Appendix I of the final published version of Wang and Shah (2024) includes a matching lower bound proof when $T$ is binary and the PS model follows a special GLM but the OR model is still linear, by adapting the proof strategy of Cai et al. (2023).
@@
The current analysis still requires that the observed data $O=(X,T,Y)$ be light-tailed. It is of great interest to investigate if the proposed approach can be extended into settings in which (part of) $O$ has heavier tails (Kuchibhotla and Chakrabortty, 2022) or has even diverging higher-order moments (e.g. transformed covariates by B-splines or Cohen-Daubechies-Vial wavelets (Belloni et al., 2015; Liu et al., 2017; Mukherjee and Sen, 2018; Liu et al., 2021)) [^7] Based on personal communications (Mukherjee, 2023), R. Mukherjee and collaborators have established nontrivial (and possibly tight) lower bounds for certain doubly-robust functionals in this context in an unpublished technical report.. ∎
Remark 6.
Last but not least, we briefly comment on the computational issue of the DCal approach at large. Depending on the link functions, the resulting optimization program for DCal can be non-convex. Smucler et al. (2019) discussed extensively the potential drawback of solving non-convex programs, such as the lack of theoretical convergence guarantees. Deriving statistical properties under computational constraint (Zadik, 2019) is a very important and potentially difficult problem that does not yet garner enough attention in the causal inference literature. But in order not to lead readers astray from the main message of the paper, we leave this problem to future work. Next in Section 5, for ATE, we propose a computationally feasible estimator that is motivated by the DCal methodology. Even though we do not have a complete theoretical proof, it nonetheless exhibits competitive performance against several popular state-of-the-art methods in a variety of simulation settings. ∎
4 Applications and extensions of the Double-Calibration approach
Thus far, the new DCal approach has been delineated using the ATE under the sparse GLM nuisance models. To demonstrate the generality of the DCal approach, we now apply and extend it in several directions.